OpenStudy (anonymous):

Parkash bought a new car at the dealership for $27,000. It is estimated that the value of the car will decrease 7% each year. Which exponential function models the value v of the car after t years?

OpenStudy (anonymous):

i got v=27,000(0.93)^t is that correct?

OpenStudy (mven):

im pretty sure it is V=27000-(0.07*t)

OpenStudy (anonymous):

no, the answers they give me are A.v=27,000(1.03)^t B.v=27,000(0.93)^t C.v=27,000(0.3)^t D.v=27,000(1.3)^t

OpenStudy (imstuck):

i got this, according to the exponential decay formula:\[V=27,000(1-.07)^{t}\]Or simplfying,\[V=27,000(.93)^{t}\]

OpenStudy (mven):

I dont know where the decimals are coming from for those sorry

OpenStudy (imstuck):

It's B. Look at my formula above.

OpenStudy (imstuck):

Value at the end = Original amount (1 - rate)^t

OpenStudy (anonymous):

yea that's what i thought too, thanks you so much!

OpenStudy (imstuck):

You're welcome!

OpenStudy (imstuck):

TY for the medal!

OpenStudy (anonymous):

You're welcome lol (: , i might be posting more questions soon hahahaha

OpenStudy (imstuck):

Tag me if you want and I will try to help you.

OpenStudy (anonymous):

Sure thing!